# MMiC: Mitigating Modality Incompleteness in Clustered Federated Learning

### v1

@@
Abstract
In the era of big data, data mining has become indispensable for uncovering hidden patterns and insights from vast and complex datasets. The integration of multimodal data sources further enhances its potential. Multimodal Federated Learning (MFL) is a distributed approach that enhances the efficiency and quality of multimodal learning, ensuring collaborative work and privacy protection. However, missing modalities pose a significant challenge in MFL, often due to data quality issues or privacy policies across the clients. In this work, we present MMiC, a framework for Mitigating Modality incompleteness in MFL within the Clusters. MMiC replaces partial parameters within client models inside clusters to mitigate the impact of missing modalities. Furthermore, it leverages the Banzhaf Power Index to optimize client selection under these conditions. Finally, MMiC employs an innovative approach to dynamically control global aggregation by utilizing Markovitz Portfolio Optimization. Extensive experiments demonstrate that MMiC consistently outperforms existing federated learning architectures in both global and personalized performance on multimodal datasets with missing modalities, confirming the effectiveness of our proposed solution.
Machine Learning, ICML
1 Introduction
Federated Learning (FL), introduced by McMahan et al. (McMahan et al., 2017), enables multiple participants to collaboratively train a unified model without sharing raw data, thus preserving privacy. This paradigm is particularly well-suited for applications such as medical image analysis, connected vehicles, blockchain and so on(Kaissis et al., 2020; Pokhrel & Choi, 2020; Lu et al., 2020). However, the absence of data sharing in FL introduces significant challenges, including model heterogeneity, communication overhead, and data distribution disparities.
@@ 1 Introduction
Clustered Federated Learning (CFL), a specialized form of FL Personalization (Tan et al., 2023), groups clients into clusters based on models, parameters, or datasets. Collaborative training within each cluster reduces the impact of data distribution disparities between clients. Recent advancements in CFL, such as one-shot clustering methods like PACFL (Vahidian et al., 2023), leverage the decomposability of data to identify client clusters efficiently. Each cluster inherently contains clients with similarities in both data and model parameters. We can leverage this characteristic of federated clustering to improve the training scenarios involving missing modalities. This means that the model parameters from clients within the same cluster that have no missing modalities can be used to fill in the parameters of clients with missing modalities. Due to the constraints of communication efficiency within clusters, we can shrink replaced parameters to only one layer most significantly affected by the missing modalities.
Current CFL algorithms predominantly employ random client selection (Ghosh et al., 2019, 2020; Ruan & Joe-Wong, 2022), which ensures fairness over multiple rounds of updates by granting each client an equal opportunity to participate in aggregation. However, we do not expect clients with severe missing modalities to enjoy the same level of aggregation rights. Therefore, by analyzing performance variations across client training rounds, we can determine whether a client is frequently affected by missing modalities. This allows us to adjust the probability of client selection accordingly.
@@
To achieve these goals, we propose MMiC framework that integrates our solutions for the three issues. In particular, we employ the rate of parameter variation to identify the parameters to be replaced. Moreover, the parameters from clients possessing complete modality data and belonging to the same cluster will be utilized to supplant those of clients with incomplete modality data. To enhance the current cluster client selection strategy in the presence of missing modalities, we incorporate the Banzhaf Power Index (BPI), a well-known coalitional game theory, to assess the contribution of each client. This approach enables us to preferentially select the clients who are more contributable, to perform more efficient and effective aggregation processes. We further employ Markovitz Portfolio Optimization, a return-risk balancing asset optimization algorithm to assesses the training process on clusters in a nonlinear manner compared to the standard weighted averaging (Wang & Ji, 2024). The global model adjusts the weights of parameter updates based on the scores obtained from the clusters, thereby reducing the impact of missing modalities in local clients on the global model.
@@
  • We are the first to apply the Markowitz Portfolio Optimization theory to evaluate the training on clusters by incorporating the evaluation results into the global aggregation process to mitigate the negative impact from missing modalities to the global server.
2 Related Works
2.1 Clustered Federated Learning (CFL)
Centralized federated learning trains a global model in a central server by aggregating gradients collected across numerous participating clients. The heterogeneity of clients in federated learning introduces challenges such as client drift, feature drift, and label drift, which significantly affect the performance of FL models. Personalized federated learning (PFL) has sprung up designed to better suit the individual characteristics of each client to address the heterogeneity. PFL strategies like Multi-Task Learning, Personalization Layers and Clustered Federated Learning (CFL) have attracted considerable attention in the literature (Smith et al., 2017; Tan et al., 2023; Mansour et al., 2020; Wu et al., 2020).
The fundamental concept of CFL is to group clients based on the similarity of their model parameters, data, or gradients. CFL emphasizes improving cooperation within these clusters to optimize the performance of cluster-specific models, while also enhancing overall global model performance. Ghosh et al. (Ghosh et al., 2019) first introduced the architecture for CFL, and recommended clustering based on the similarity of model parameters. This architecture has been refined through iterative clustering techniques such as the Iterative Federated Clustering Algorithm (IFCA) (Sattler et al., 2021). Fedsoft further advanced the iterative CFL by allowing a client to belong to multiple clusters, outperforming previous efforts. However, these methods incurred substantial resource consumption. One-shot clustering was then introduced by Dennis et al. (Dennis et al., 2021) for more efficient clustering. Their proposed method efficiently determined clustering attribution, identifying the uniqueness of client metrics within complex federated architectures. Vahidian et al. (Vahidian et al., 2023) proposed a one-shot CFL algorithm based on data subspace clustering, They measured the similarity between clients by decomposing the dataset via Singular Value Decomposition (SVD). Our proposed MMiC framework is under the same setting as the one-shot CFL (Vahidian et al., 2023; Yang et al., 2024) and extends it to multimodal scenarios with missing modalities.
2.2 Multimodal Federated Learning (MFL)
@@
Challenges and Advances in MFL. FedCMR is the first to address multimodal tasks, though it encountered challenges such as modality discrepancies and missing data. Recent works have made strides in tackling these issues. For instance, Chen and Li (Chen & Li, 2022) proposed MFL-HGB, which adaptively computes optimal modality blending and local model weighting to mitigate overfitting. Similarly, Fed-MP employs adaptive aggregation of local model weights and a multimodal prototyping mechanism to enhance generalization to unseen classes (Zeng et al., 2024). These works collectively provide a foundation for understanding the challenges and solutions in MFL (Che et al., 2023).
@@
MFL with Missing Modality. CACMRN (Xiong et al., 2023) is the first FL framework to propose solutions for missing modalities. They aim to achieve modality reconstruction using existing multi-modal data by training multiple functionally diverse models. Yu et al. (Yu et al., 2023) aim to address the issue of missing modalities in cross-modal retrieval. Clients, each with distinct modalities, learn the representation information of specific modalities and upload it to the server for contrastive learning. Fed-Multimodal (Feng et al., 2023) is a multimodal benchmark built upon the FL algorithm. The work presented a range of models and datasets and proposed a method to initialize the missing modal vector as zero. PmcmFL (Bao et al., 2024) applied the common prototype exchange method of federated learning against Non-IID to the missing modality, and they store prototypes of all modalities for all classes from each client on the server. When encountering missing modalities, it utilizes the corresponding class-specific prototypes to fill in the representations. Additionally, they applied the MIX-UP (Zhang et al., 2018) method in data augmentation into the additional training dataset. In contrast, our approach does not impose restrictions on the types of multimodal tasks, meaning that we do not require the use of specific models.
3 Preliminaries
Before jumping into the technical details of MMiC, we first introduce the fundamental concepts of CFL and the relevant concepts in the game theory and finance domain that have been adopted by MMiC. To assist the description, we list the notations used throughout the paper in Table 1.
Table 1: Main notations.
Notation Description
$K$ Number of clients
$c$ Client
$M$,$m$ Number, index of cluster
$\mathcal{G},\mathcal{M}$ Server, cluster
$T$, $t$ Global Training rounds
$E$, $e$ Local training epochs
$B$, $b$ Batch number, b-th batch
$I,i$, $J,j$ Temporary index or sequence
$\Theta,\theta$ Model or parameters
$\ell$, $L$ Local loss, global Loss
$\eta$ Learning rate
$g$ Gradient
$\mathcal{D}$ Dataset
$R$ Client joining rate each round
$\gamma_{t}$ Portfolio score in $t$-th round
$a,\alpha$ Performance gap
$\sigma$ Risk of federated cluster
$\phi(*)$ Banzhaf interaction index
$w_{k}$ Weight of client $k$
$P$, $\mathcal{P}$ Number of layers, the poverty layer
$\lambda,\epsilon,\tau$ Control constant
$\mathcal{S}$ Subset
3.1 Clustered Federated Learning
Given a global server $\mathcal{G}$ and $K$ clients $\{c_{i}\}^{K}_{1}$, let $\mathcal{D}_{k}$ over $\mathcal{X}*\mathcal{Y}$ be the local dataset of the $k$-th client $c_{k}$, with $\mathcal{X}$ and $\mathcal{Y}$ represent the data inputs and labels, $n_{k}=|\mathcal{D}_{k}|$ be the dataset size. Client weight is calculated as $w_{k}=\frac{n_{k}}{\sum_{k}^{K}n_{k}}$. $c_{k}$ and $\mathcal{G}$ possess the model $\theta_{k}$ and $\theta_{\mathcal{G}}$ respectively and the models are the same when initialized.
Learning Objective. During the FL training process, each cluster model, on the one hand, is regarded as a mesosphere agent that can optimize the training loss from clients within the cluster; and on the other hand, it serves as a local model to the global model, providing stronger generalization ability. In each round of federated training, on the client, it is to learn local model $\theta_{k}$ by minimizing the local empirical loss:
$F_{i}(\theta_{k})=\sum_{\xi\in\mathcal{D}_{k}}\frac{\ell(\theta_{k},\xi)}{n_{k}}$ (1)
@@ 3.1 Clustered Federated Learning
where $\ell(\cdot,\cdot)$ is the pre-defined task-driven loss function and $\xi$ represents sampled training data in a batch. The clusters are defined by partitioning $K$ clients into $M$ disjoint subsets. Different from traditional federated selection, clustered federated selection is refined to each cluster, and each round has client subset $\mathcal{S}_{m}$ selected in $\mathcal{M}_{m}$. The weight of cluster $m$ is usually calculated as $w_{m}=\frac{\sum_{k\in\mathcal{M}_{m}}n_{k}}{\sum_{1}^{M}\sum_{k\in\mathcal{M}_{m}}n_{k}}$. Therefore, the global objective function can be expressed as:
$\quad min\quad\sum_{m=1}^{M}\sum_{k\in\mathcal{M}_{m}}w_{k}\ell(\boldsymbol{\theta}_{k})$ (2)
@@
Model Aggregation. A standard way to obtain a cluster model from clients’ models in CFL is to adopt the FedAvg (Li et al., 2020):
$\boldsymbol{\Theta}_{m}=\sum_{k\in\mathcal{M}_{m}}w_{k}\boldsymbol{\theta}_{k}$ (3)
There are various aggregation methods applied to aggregate the global model from clusters’ models. FedAdagrad (Reddi et al., 2021b) exploits dynamic aggregation to combine the clients’ parameter updates from the current round with the stored historical parameter updates:
$\Delta_{m,t}=\boldsymbol{\Theta}_{m,t}-\boldsymbol{\Theta}_{\mathcal{G},t-1}$ (4)
$\boldsymbol{\Delta}_{t}=\frac{1}{M}\sum_{m=1}^{M}w_{m}\boldsymbol{\Delta}_{m,t}$
$\boldsymbol{mtumn}_{t}=\beta\boldsymbol{mtumn}_{t-1}+(1-\beta)\Delta_{t},\beta\in[0,1)$
$\boldsymbol{\upsilon}_{t}=\boldsymbol{\upsilon}_{t-1}+\boldsymbol{\Delta}_{t}^{2}$
$\boldsymbol{\theta}_{\mathcal{G},t}=\boldsymbol{\theta}_{\mathcal{G},t-1}+\eta\frac{\boldsymbol{mtumn}_{t}}{\sqrt{\boldsymbol{\upsilon}_{t}}+\tau}$
where $\Delta_{m,t}$ [^1] Both $\Theta_{m,t}$ and $\Theta_{m}$ represent the cluster model. Here we add the time stamp $t$ to denote $t$-th round of updating and $\boldsymbol{\Delta}_{t}$ is the difference between the current cluster model and the global model. $\boldsymbol{mtumn}_{t}$ is updated iteratively and collects the historical parameter updates, while $\beta$ balances the historical parameter updates $\boldsymbol{mtumn}_{t-1}$ and $\boldsymbol{\Delta}_{t}$. A larger $\beta$ indicates the retention of more historical parameter information. $\boldsymbol{\upsilon}_{t}$ is an accumulated squared change estimate used to measure changes in parameters and adjust the learning rate to improve the stability and convergence speed of the optimization process. MMiC relies on FedAdagrad and derives an extended version (Section 4.3).
3.2 Banzhaf Power Index
The Banzhaf Power Index (BPI) is a game theory metric used to quantify the power of an individual voter in a decision-making body where votes are not necessarily equal (Dubey & Shapley, 1979; Laruelle & Valenciano, 2001). The index is calculated by considering all possible coalitions (i.e., combinations of voters) that can be formed within a voting body and determining the importance of each voter in each coalition. A voter $k$ is considered a core member if their presence in a coalition changes the outcome (denoted as $A_{\mathcal{S}}$) from a loss to a win:
$k\,is\,core?\begin{cases}True&\text{if }A_{\mathcal{S}^{\prime}}\,is\,a\,loss\text{ and }A_{\mathcal{S}}\,is\,a\,win\\ False&\text{otherwise}\end{cases}$ (5)
where $\mathcal{S}^{\prime}$ is the set that excludes voter $k$: $\mathcal{S}^{\prime}=\mathcal{S}\setminus\{k\}$. BPI could be adopted in federated learning settings to help select the most contributable clients (i.e., a subset of the coalition) during the training process, making the client selection more effective.
3.3 Markowitz Portfolio Optimzation
The Markowitz Portfolio Optimization model is an investment optimization model that aims to obtain an optimal investment portfolio with maximized returns and minimized risks (Miller, 1960). The model can assess the current investment process through Risk-Adjusted Return Coefficient RARC ($\gamma$). The quadratic programming model of Markowitz Portfolio Optimization is formulated as follows:
$minimize\quad\gamma=\lambda\bar{\sigma}-(1-\lambda)A$ (6)
indicating RARC is influenced by three variables: investment risk $\bar{\sigma}$, the investment return $A$ and the risk tolerance factor $\lambda$. $\lambda$ signifies the investor’s risk tolerance within the range of [0,1] - When $\lambda$ is set to 0.0, the maximum return value is achieved. $A$ denotes the total portfolio return across all assets. $\bar{\sigma}$ is calculated as follows:
$\bar{\sigma}=\sum_{i=1}^{K}\sum_{j=1}^{K}w_{i}w_{j}\sigma_{ij}$ (7)
$\sigma_{ij}=\frac{1}{T-1}\sum_{t=1}^{T}(\alpha_{i,t}-\bar{\alpha}_{i,t-1})(\alpha_{j,t}-\bar{\alpha}_{j,t-1})$
where $\sigma_{ij}$ is the covariance between assets $i$ and $j$. $w_{i}$ and $w_{j}$ are the weights of assets $i$ and $j$, respectively. $K$ is the number of assets. $\bar{\alpha}_{i,t-1}$ denotes the historical average return of assets $i$ across the preceding $t-1$ rounds. $\sigma_{ij}$, the risk between assets $i$ and $j$ is obtained by subtracting the corresponding average return from each round in the history of the asset. This formulation can be extended to federated learning scenarios, where $\sigma_{ij}$ quantifies the risk or divergence of local models, and $\bar{\alpha_{i,t-1}}$ reflect the historical performance of individual participants in the federated system.
4 Methodology

### v3

@@
Abstract.
In the era of big data, data mining has become indispensable for uncovering hidden patterns and insights from vast and complex datasets. The integration of multimodal data sources further enhances its potential. Multimodal Federated Learning (MFL) is a distributed approach that enhances the efficiency and quality of multimodal learning, ensuring collaborative work and privacy protection. However, missing modalities pose a significant challenge in MFL, often due to data quality issues or privacy policies across the clients. In this work, we present MMiC, a framework for Mitigating Modality incompleteness in MFL within the Clusters. MMiC replaces partial parameters within client models inside clusters to mitigate the impact of missing modalities. Furthermore, it leverages the Banzhaf Power Index to optimize client selection under these conditions. Finally, MMiC employs an innovative approach to dynamically control global aggregation by utilizing Markowitz Portfolio Optimization. Extensive experiments demonstrate that MMiC consistently outperforms existing federated learning architectures in both global and personalized performance on multimodal datasets with missing modalities, confirming the effectiveness of our proposed solution. Our code released at https://github.com/gotobcn8/MMiC.
Federated Learning, Multimodal Learning, Missing Modality
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761140ccs: Computing methodologies Distributed computing methodologiesccs: Information systems Data miningccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Machine learning
1. Introduction
Federated Learning (FL), introduced by McMahan et al. (McMahan et al., 2017), enables multiple participants to collaboratively train a unified model without sharing raw data, thus preserving privacy. This paradigm is particularly well-suited for applications such as medical image analysis, connected vehicles, blockchain and so on(Kaissis et al., 2020; Pokhrel and Choi, 2020; Lu et al., 2020). However, the absence of data sharing in FL introduces significant challenges, including model heterogeneity, communication overhead, and data distribution disparities.
@@ 1. Introduction
Clustered Federated Learning (CFL), a specialized form of FL Personalization (Tan et al., 2023), groups clients into clusters based on models, parameters, or datasets. Collaborative training within each cluster reduces the impact of data distribution disparities between clients. Recent advancements in CFL, such as one-shot clustering methods like PACFL (Vahidian et al., 2023), leverage the decomposability of data to identify client clusters efficiently. Each cluster inherently contains clients with similarities in both data and model parameters. We can leverage this characteristic of federated clustering to improve the training scenarios involving missing modalities. This means that the model parameters from clients within the same cluster that have full modalities can be used to fill in the parameters of clients with missing modalities. Due to the constraints of communication efficiency within clusters, we can shrink replaced parameters to only one layer most significantly affected by the missing modalities.
Current CFL algorithms predominantly employ random client selection (Ghosh et al., 2019, 2020; Ruan and Joe-Wong, 2022), which ensures fairness over multiple rounds of updates by granting each client an equal opportunity to participate in aggregation. However, we do not expect clients with severe missing modalities to enjoy the same level of aggregation rights. Therefore, by analyzing performance variations across client training rounds, we can determine whether a client is frequently affected by missing modalities. This allows us to adjust the probability of client selection accordingly.
@@
To achieve these goals, we propose MMiC, a framework that integrates our solutions for the three issues. In particular, we employ the rate of parameter variation to identify the parameters to be replaced. Moreover, the parameters from clients possessing complete modality data and belonging to the same cluster will be utilized to supplant those of clients with incomplete modality data. To enhance the current cluster client selection strategy in the presence of missing modalities, we incorporate the Banzhaf Power Index (BPI), a well-known coalitional game theory, to assess the contribution of each client. This approach enables us to preferentially select the clients who are more contributable, to perform more efficient and effective aggregation processes. We further make use of Markowitz Portfolio Optimization, a return-risk balancing asset optimization algorithm to assess the training process on clusters in a nonlinear manner compared to the standard weighted averaging (Wang and Ji, 2024). The global model adjusts the weights of parameter updates based on the scores obtained from the clusters, thereby reducing the impact of missing modalities in local clients on the global model.
@@
  • We are the first to apply the Markowitz Portfolio Optimization theory to evaluate the training process by incorporating the evaluation results into the global aggregation process to mitigate the negative impact of missing modalities on the global server.
2. Related Works
2.1. Clustered Federated Learning (CFL)
Centralized federated learning trains a global model in a central server by aggregating gradients collected across numerous participating clients. Personalized federated learning (PFL) has sprung up designed to better suit the individual characteristics of each client to address the heterogeneity. PFL strategies like Multi-Task Learning, Personalization Layers and Clustered Federated Learning (CFL) have attracted considerable attention in the literature (Smith et al., 2017; Tan et al., 2023; Mansour et al., 2020; Wu et al., 2020).
The fundamental concept of CFL is to group clients based on the similarity of their model parameters, data, or gradients. CFL emphasizes improving cooperation within these clusters to optimize the performance of cluster-specific models, while also enhancing overall global model performance. Ghosh et al. (Ghosh et al., 2019) first introduced the architecture for CFL, and recommended clustering based on the similarity of model parameters. Fedsoft further advanced the iterative CFL by allowing a client to belong to multiple clusters, outperforming previous efforts. However, these methods incurred substantial resource consumption. One-shot clustering was then introduced by Dennis et al. (Dennis et al., 2021) for more efficient clustering. Their proposed method efficiently determined clustering attribution, identifying the uniqueness of client metrics within complex federated architectures. Vahidian et al. (Vahidian et al., 2023) proposed a one-shot CFL algorithm based on data subspace clustering, They measured the similarity between clients by decomposing the dataset via Singular Value Decomposition (SVD). Our proposed MMiC framework is under the same setting as the one-shot CFL (Vahidian et al., 2023; Yang et al., 2024) and extends it to multimodal scenarios with missing modalities.
2.2. Multimodal Federated Learning (MFL)
@@
Challenges and Advances in MFL. FedCMR is the first to address multimodal tasks, though it encountered challenges such as modality discrepancies and missing data. Recent works have made strides in tackling these issues (Zong et al., 2021). Graph-based FL represents a special case of CFL, they could be used to handle the heterogeneous modality fusion. FedMSplit (Chen and Zhang, 2022) employs a dynamic and multi-view graph structure to adaptively capture the correlations amongst multimodal local models. AGHG (Qi and Li, 2024) creates a global knowledge enhancer to extract information from different modalities. These works collectively provide a foundation for understanding the challenges and solutions in MFL (Che et al., 2023).
@@
MFL with Missing Modality. CACMRN (Xiong et al., 2023) is the first FL framework to propose solutions for missing modalities. They aim to achieve modality reconstruction using existing multi-modal data by training multiple functionally diverse models. Yu et al. (Yu et al., 2023) aim to address the issue of missing modalities in cross-modal retrieval. Clients, each with distinct modalities, learn the representation information of specific modalities and upload it to the server for contrastive learning. Fed-Multimodal (Feng et al., 2023) is a multimodal benchmark built upon the FL algorithm. The work presented a range of models and datasets and proposed a method to initialize the missing modal vector as zero. PmcmFL (Bao et al., 2024) applied the common prototype exchange method of federated learning against Non-IID to the missing modality, and they store prototypes of all modalities for all classes from each client on the server. When encountering missing modalities, it utilizes the corresponding class-specific prototypes to fill in the representations. In contrast, our approach does not impose restrictions on the types of multimodal tasks, meaning that we do not require the use of specific models.
3. Clustered Federated Learning
@@
Optimization Objective. Given a global server $\mathcal{G}$ and $K$ clients $\{c_{i}\}^{K}_{1}$, let $\mathcal{D}_{k}$ over $\mathcal{X}*\mathcal{Y}$ be the local dataset of the $k$-th client $c_{k}$, with $\mathcal{X}$ and $\mathcal{Y}$ represent the data inputs and labels, $n_{k}=|\mathcal{D}_{k}|$ be the dataset size. Client aggregated weight is calculated as $w_{k}=\frac{n_{k}}{\sum_{k}^{K}n_{k}}$. $c_{k}$ and $\mathcal{G}$ possess the model $\theta_{k}$ and $\theta_{\mathcal{G}}$ respectively and the models are the same when initialized. In each round of federated training, on the client, it is to learn local model $\theta_{k}$ by minimizing the local empirical loss: $F_{i}(\theta_{k})=\sum_{\xi\in\mathcal{D}_{k}}\frac{\ell(\theta_{k},\xi)}{n_{k}}$. Where $\ell(\cdot,\cdot)$ is the pre-defined task-driven loss function and $\xi$ represents sampled training data in a batch. The clusters are defined by partitioning $K$ clients into $M$ disjoint subsets. Different from traditional federated selection, clustered federated selection is refined to each cluster, and each round has client subset $\mathcal{S}_{m}$ selected in $\mathcal{M}_{m}$. The weight of cluster $m$ is usually calculated as $w_{m}=\frac{\sum_{k\in\mathcal{M}_{m}}n_{k}}{\sum_{1}^{M}\sum_{k\in\mathcal{M}_{m}}n_{k}}$. Therefore, the global objective function can be expressed as:
(1) $\quad min\quad\sum_{m=1}^{M}\sum_{k\in\mathcal{M}_{m}}w_{k}\ell(\bm{\theta}_{k})$
@@
Model Aggregation. We obtain cluster models from clients’ models by using FedAvg (Li et al., 2020): $\bm{\Theta}_{m}=\sum_{k\in\mathcal{M}_{m}}w_{k}\bm{\theta}_{k}$. Our method for aggregating cluster models into the global model is elaborated in Section 4.3.
4. Methodology